I'm some problems with my algebra lesson. It's systems of equations. I am suppose to figure out if the ordered pair of (8,-1) is a solution to the systems: 2x + 3y = 13 and x - y = 19. I know how to do this, just not with this problem. On the second equation I am confused as to what the x and the y are supposed to be, number wise. What do I do with the x and y when there is no other number?
Answer by Alan3354(69443)

I am suppose to figure out if the ordered pair of (8,-1) is a solution to the systems: 2x + 3y = 13 and x - y = 19. I know how to do this, just not with this problem. On the second equation I am confused as to what the x and the y are supposed to be, number wise. What do I do with the x and y when there is no other number?
-----------------
Sub 8 for x and -1 for y, and see if the equations are true.
2x + 3y = 13
2*8 + 3(-1) =? 13
16 - 3 = 13 That one is true.
------------------
x - y = 19
8 - (-1) =? 19
9 <> 19 so it's not a solution to the 2nd equation.
If you graph these, the point (8,-1) is on the line of the 1st eqn, but not on the 2nd.